OpenStudy (anonymous):

The perimeter of a rectangle is 120 ft. the length is five times the width. What is the length and width?

OpenStudy (anonymous):

2l+2w=120, l=5*w

OpenStudy (anonymous):

So 5*w+w=60 => w = 10 and l = 50.